A BALL IS RELEASED FROM HIGHT AND REBOUNDS WITH SAME SPEED AFTER HITTING THE GROUNDIF IT TRAVELS 20M IN FIRST 8 SECONDS FIND THE NOOF COLLISIONS MADE)
1
Expert's answer
2020-09-06T17:15:55-0400

"(2N-1)t=8" and "(2N-1)\frac{gt^2}{2}=20"


We have


"\frac{40}{2N-1}=g\frac{64}{(2N-1)^2} \to N=\frac{64}{80}g+0.5="


"=\frac{64}{80}\cdot9.81+0.5 \approx8."


So, "N=8".